OpenStudy (volleyballlover55):

The distance of planet Saturn from the Sun is 1.5 ⋅ 10^9 kilometers, and the distance of planet Venus from the Sun is 1.1 ⋅ 10^8 kilometers. About how many more kilometers is the distance of Saturn from the Sun than the distance of Venus from the Sun? 0.4 ⋅ 10^8 0.4 ⋅ 10^9 1.4 ⋅ 10^9 1.4 ⋅ 10^10

OpenStudy (volleyballlover55):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

what is 10^9-10^8?

OpenStudy (volleyballlover55):

10^1

rebeccaxhawaii (rebeccaxhawaii):

wait hold on i need to refresh my brain do you have a formula to subtract exponents? @rishavraj

OpenStudy (volleyballlover55):

are you refreshed? @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

@4everaddicted2anime

OpenStudy (volleyballlover55):

??

OpenStudy (4everaddicted2anime):

I am a little confused. If we are finding the difference between the two then we would subtract the number with exponents and then separately subtract the numbers in the beginning. I know that 10^9 - 10^8 is 10^1. But that isn't in one of the possible answers?

OpenStudy (volleyballlover55):

i know thats what i thought

OpenStudy (volleyballlover55):

@Bobo-i-bo

OpenStudy (bobo-i-bo):

Hint: \[1.1\times10^8 = 1.1\times10^8 \times \frac{10}{10} = \frac{1.1}{10}(10^8 \times 10) = 0.11 \times 10^9\]